Draw a rectangle that has
A= 50 sq cm
P=54 cm
Do i just guess and check or how do I find the answer?
THANKS

Answer by richard1234(7193)

If you let L and W be the length and width then we have

LW = 50

2L + 2W = 54 --> L + W = 27

You could guess and check to obtain L = 25, W = 2. The surefire method is to let L = 27 - W, then substitute into the first equation to get (27-W)(W) = 50, then solve for W and L.